# The Curious Case of In-Training Compression of State Space Models

### v2

@@ A.1 Solving The Lyapunov Equations for Diagonal SSMs
For SSMs with diagonal state transition matrix ${\bm{A}}=\text{diag}(a_{1},\ldots,a_{n})$, which covers a lot of SSMs used today such as LRU Orvieto et al. (2023) and S5 Smith et al. (2022), Equation 2 and Equation 3 admit a simple, entry-wise closed-form solution:
${\bm{P}}_{ij}$ $=\frac{\left({\bm{B}}{\bm{B}}^{\top}\right)_{ij}}{1-a_{i}a_{j}},\quad{\bm{Q}}_{ij}=\frac{\left({\bm{C}}{\bm{C}}^{\top}\right)_{ij}}{1-a_{i}a_{j}}\quad\forall 1\leq i,j\leq n$ (14)
@@
A.2 Selective SSM case
Selective SSMs are built with dynamical systems that fall under the linear parameter varying (LPV) framework, with the parameter being the layer input. Such systems are also referred to as Linear Input Varying (LIV) and their general case state equations are given by,
${\bm{y}}(k)$ $={\bm{C}}({\bm{x}}(k))\,{\bm{h}}(k)+{\bm{D}}({\bm{x}}(k))\,{\bm{x}}(k),$ (16)
so that the Gramians $\bm{P}$} and $\bm{Q}$} are input-dependent.
The caveat is that the mean system may not be stable, controllable, or observable; one may therefore need to regularize the mean matrices to satisfy these assumptions before applying a single global reduction based on the LTI approach.
Appendix B Experimental Details
@@
Table 2: Hyperparameters for LRU experiments. “h“ refers to the dimension of the hidden state, “n“ to the state space dimension, “Batch“ to the batch size and “LR Factor“ to the learning rate factor applied to the sequence mixer Orvieto et al. (2023).
Task Depth h n Steps Batch LR Factor Weight Decay Dropout
MNIST 11 88 $256$ $200k$ $50$ - - $0.1$
@@
On LRA tasks, the learning rate is warmed up from $10^{-7}$ to $10^{-3}$ for $10\%$ of the total steps, before it is cosine-decayed back to $10^{-7}$. For MNIST, the learning rate is fixed at $4\cdot 10^{-4}$ for the entirety of training.
B.1 Reduction Details
We find that LRU overfits on IMDB, even after reducing the number of parameters by 6 and doubling the dropout rate compared to Orvieto et al. (2023). Training LRU on this dataset, we furthermore observe an initial training period in which the loss plateaus. Just on IMDB, we thus wait for an initial $1k$ steps before doing the balanced truncation. Instead of doing 4 reduction steps until the end of warmup, we also just do 2 until $3k$ steps in order to give the model time to recover before entering the overfitting regime.
Appendix C Additional Results
C.1 Performance
In Figure 4 we provide the state dimension vs test performance plots for all datasets.
(a) MNIST

### v3

@@ A.1 Solving The Lyapunov Equations for Diagonal SSMs
For SSMs with diagonal state transition matrix ${\bm{A}}=\text{diag}(\lambda_{1},\ldots,\lambda_{n})$, which covers a lot of SSMs used today such as LRU (Orvieto et al., 2023) and S5 (Smith et al., 2022), Equation 2 and Equation 3 admit a simple, entry-wise closed-form solution:
${\bm{P}}_{ij}$ $=\frac{\left({\bm{B}}{\bm{B}}^{\top}\right)_{ij}}{1-\lambda_{i}\lambda_{j}},\quad{\bm{Q}}_{ij}=\frac{\left({\bm{C}}^{\top}{\bm{C}}\right)_{ij}}{1-\lambda_{i}\lambda_{j}}\quad\forall 1\leq i,j\leq n$ (14)
@@
A.2 Hyperparameter Selection
@@
Table 3: Hyperparameters for LRU experiments. $h$ refers to the dimension of the hidden state, $n$ to the state space dimension, $B$ to the batch size and $\alpha_{LR}$ to the learning rate factor applied to the sequence mixer (Orvieto et al., 2023).
Task Depth $\bm{h}$ $\bm{n}$ Steps $\bm{B}$ $\bm{\alpha_{LR}}$ Weight Decay Dropout
sMNIST 11 88 $256$ $200k$ $50$ - - $0.1$
@@
On LRA tasks, the learning rate is warmed up from $10^{-7}$ to $10^{-3}$ for $10\%$ of the total steps, before it is cosine-decayed back to $10^{-7}$. For sMNIST, the learning rate is fixed at $4\cdot 10^{-4}$ for the entirety of training.
A.3 Reduction Details
We find that LRU overfits on IMDB, even after reducing the number of parameters by 6 and doubling the dropout rate compared to Orvieto et al. (2023). Training LRU on this dataset, we furthermore observe an initial training period in which the loss plateaus. Just on IMDB, we thus wait for an initial $1k$ steps before doing the balanced truncation. Instead of doing 4 reduction steps until the end of warmup, we also just do 2 until $3k$ steps in order to avoid entering the overfitting regime.
Appendix B Additional Numerical Results
This appendix expands upon the empirical findings presented in the main paper by providing a more complete set of numerical results and supporting analyses. We aim to complete the view of how CompreSSM behaves, and to substantiate the assumptions underlying our in-training reduction framework. We first report extended performance curves and dimension statistics across all benchmarks, results for the pragmatic variant of the algorithm, followed by additional evaluations of the dynamical behavior of Hankel singular values throughout training. Together, these results offer a holistic understanding of the stability, robustness, and practical effectiveness of the proposed method.
B.1 Performance
In Figure 4 we provide the state dimension vs test performance plots for all datasets. These follow the same conventions and experimental setup as those described in the main text for Subfigure (3(a)). The baselines again correspond to models that are trained from scratch at the with the mean of the reduced dimensions reached at the various tolerance levels. Time gain plots are not provided, but similar gains are observed across all experiments, with an in depth discussion and analysis on the subject provided in Section D of this appendix.
(a) sMNIST